# An Approach for Auto Generation of Labeling Functions for Software Engineering Chatbots

### v1

@@
Abstract
Software engineering (SE) chatbots are increasingly gaining attention for their role in enhancing development processes. At the core of chatbots are the Natural Language Understanding platforms (NLUs), which enable them to comprehend and respond to user queries. Before deploying NLUs, there is a need to train them with labeled data. However, acquiring such labeled data for SE chatbots is challenging due to the scarcity of high-quality datasets. This challenge arises because training SE chatbots requires specialized vocabulary and phrases not found in typical language datasets. Consequently, chatbot developers often resort to manually annotating user queries to gather the data necessary for training effective chatbots, a process that is both time-consuming and resource-intensive. Previous studies propose approaches to support chatbot practitioners in annotating users’ posed queries. However, these approaches require human intervention to generate rules, called labeling functions (LFs), that identify and categorize user queries based on specific patterns in the data. To address this issue, we propose an approach to automatically generate LFs by extracting patterns from labeled user queries. We evaluate the effectiveness of our approach by applying it to the queries of four diverse SE datasets (namely AskGit, MSA, Ask Ubuntu, and Stack Overflow) and measure the performance improvement gained from training the NLU on the queries labeled by the generated LFs. We find that the generated LFs effectively label data with AUC scores of up to 85.3%, and NLU’s performance improvement of up to 27.2% across the studied datasets. Furthermore, our results show that the number of LFs used to generate LFs affects the labeling performance. We believe that our approach can save time and resources in labeling users’ queries, allowing practitioners to focus on core chatbot functionalities rather than on manually labeling queries.
Index Terms:
@@ Abstract
1 Introduction
@@
In the field of software engineering (SE), chatbots are deployed as conversational tools to automate a wide range of tasks, from assisting in problem-solving to answering questions about repositories [1, 2, 3]. At the core of these chatbots are Natural Language Understanding platforms (NLUs). These NLUs serve as the backbone of the chatbot, responsible for interpreting human language into structured data that chatbots can act on [4]. They accomplish this by using machine learning and natural language processing techniques to dissect user queries, identifying the user’s intent and extracting key entities like bug IDs or version numbers [5]. By doing so, NLUs make it possible for the chatbot to generate responses that are contextually relevant and specific to the user’s queries.
@@
The effectiveness of NLUs heavily relies on the availability of a large volume of high-quality training data [6, 7, 8]. However, prior work shows that obtaining such data is expensive, especially in specialized domains like SE [9, 10, 11], where the domain-specific language and terminology pose unique challenges for data collection and labeling [4]. For example, common words like ‘push’, ‘fork’, and ‘commit’ have distinct meanings in the context of software development, which differ from their everyday usage [4]. This makes it difficult to rely on general-purpose datasets for training SE chatbots. Furthermore, there is a notable scarcity of publicly available, high-quality datasets for training chatbots in the SE domain [12]. As a result, gathering sufficient and relevant training data for SE chatbots often requires significant effort and resources.
@@
To obtain the necessary training data for SE chatbots, practitioners often resort to mining chatbot user queries as a data source [13]. However, chatbot developers need to label these queries before they can be used for training the NLU, introducing significant challenges [14]. First, manual labeling is labor-intensive and can incur substantial costs [15], particularly when domain expertise is essential for ensuring the accuracy of labels [16, 17]. This increases the financial burden and extends the time required to prepare data for training purposes [18, 15, 19]. Second, although semi-automated labeling methods exist, they still require human intervention, such as domain experts developing heuristics for labeling user queries [20, 21, 22]. These heuristics are also called Labeling Functions (referred hereafter simply as LFs). Specifically, LFs are programmatic rules or functions that assign labels to data points based on certain conditions or patterns. For example, an LF for an SE chatbot might label a query as ‘bug-related’ if it contains keywords like ‘error’, ‘issue’ or ‘fix’.
@@
In response to the challenges outlined above regarding data labeling, we introduce our approach that automates the process of generating LFs specifically for SE datasets. To accomplish this, our approach takes a small set of labeled data as an input and automatically analyzes and extracts patterns from it, and uses those patterns to generate the LFs capable of auto-labeling data. After the generation of LFs, we use them to label data and train the NLU of the chatbot on the trained data [23]. Our approach is organized into three main components. First, the Grouper is responsible for expanding the initial labeled data by identifying similar queries [24]. Second, the LF Generator takes on the role of extracting patterns from this expanded data to create LFs [25]. Finally, the Pruner filters out low-quality LFs out of the pool of generated LFs [26, 27]. We evaluate our approach based on four datasets (namely, AskGit, MSA, Ask Ubuntu, and Stack Overflow) used to develop chatbots for performing various SE tasks. Specifically, we aim to answer the following research questions in this paper:
@@
    How well do the generated LFs label data? We evaluate the quality of the generated LFs in terms of their effectiveness in labeling data, and subsequently, in training SE chatbots [28, 29]. Our analysis shows that the generated LFs effectively label data with AUC scores of above 75.5% for three out of the four studied datasets (except Stack Overflow). Additionally, we observe that for these three datasets, using the auto-labeled data for training can enhance the NLU’s performance, with AUC score improvements of up to 27.2%.
@@
  • We make our dataset and code publicly available to facilitate future research in this area [30].
Paper Organization. The remainder of the paper is organized as follows. Section 2 provides the background that forms the basis for our study. Section 3 presents the details of our approach and its key components. Section 4 describes the setup for our empirical study followed by Section 5 which presents the results of our RQs. Section 6 provides additional analysis on the impact of the number of LFs generated by our approach on performance and Section 7 discusses threats to the validity of our study. Finally, Section 8 reviews the related work and Section 9 concludes the paper.
2 Background
Before proceeding to our approach, we explain in this section the chatbot related terminologies, data labeling process, and LFs. Also, we briefly discuss the role of NLU-based chatbots in the era of LLMs.
2.1 Chatbot Training and Data Challenges
Software chatbots serve as the conduit between users and services [31]. Users input their queries to the chatbot in natural language, which then performs the requested action (e.g., querying the database) and responds to the user’s question. NLU platforms are the backbone for chatbots to understand the user’s question. Specifically, NLU extracts two key aspects from the input query: the topic of the user’s query or what is being asked about (known as the ‘Entity’) and what the user is asking it to do (known as the ‘Intent’).
Similar to any machine learning model, NLUs need to be trained to extract intents and entities. In particular, for each intent, the NLU needs to be trained on a set of queries that represent different ways a user could express that intent. Therefore, the training data should include a wide range of examples for each intent to capture the variety of ways in which people express their needs. Figure 1 presents a snapshot of training data for the FileCreator and IssueClosingDate intents with their training examples. Chatbot practitioners need to brainstorm the different ways that the user could ask the question for a specific intent [11] to train the NLU. Nevertheless, this is a resource-intensive and time-consuming task [13].
@@ 2.1 Chatbot Training and Data Challenges
Figure 1: Example dataset for training an SE chatbot.
Alternatively, chatbot practitioners leverage user-chatbot conversations to augment their dataset [13]. To demonstrate this, we present the user-chatbot interaction example shown in Figure 2. In the example, the user (highlighted in green) asks the chatbot software repository-related questions. For the first query, the user asks, ‘I want to know who created map.json’. Here, the intent is to find out the developer who created the file. The chatbot’s NLU analyzes this query, extracting the intent (FileCreator with a confidence score of 0.87) and the relevant entity (file_name is map.json). Based on this understanding, the chatbot then provides a response identifying the file creator.
Figure 2: Example of a user’s interaction with an SE chatbot.
@@ 2.2 Data Labeling
The process of data labeling involves assigning appropriate labels to a dataset [27, 20]. There are several data labeling techniques that are traditionally used [32, 33]. These include manual labeling, crowd-sourcing, and hiring domain experts [32, 15, 16, 34].
@@
Manual labeling is the annotation of unlabeled data by humans [14]. The process typically begins with gathering the unlabeled data [35] and reviewing them. Next, appropriate labels for intents and entities are assigned to the unlabeled data, categorizing them into classes [35]. Once the data is suitably annotated, it is incorporated into the chatbot’s training dataset to further enhance its accuracy [35]. Because of all these steps, the manual labeling process requires significant effort and is time-consuming [15]. It can also lead to inconsistencies due to human error, making it less practical for large-scale projects [8].
@@
Crowdsourcing involves distributing the task of labeling data to a large group of people, often through an online platform. This method can significantly accelerate the data labeling process by leveraging the crowd’s collective effort. Although crowdsourcing can help label a large volume of data quickly, it may not always guarantee the quality necessary for specialized domains like SE. The lack of domain-specific knowledge among the crowd can lead to inaccuracies in labeling, affecting the overall quality of the training data [17, 36, 37].
@@
Recruiting domain experts refers to engaging individuals with specialized knowledge in a particular area, such as SE, to label the data. These experts bring a high level of accuracy and insight to the labeling process, ensuring the data is correctly annotated with the appropriate intents and entities. However, this method can be costly and may not scale well for large datasets, making it a challenging option for projects with limited budgets [16, 17]. Due to the limitations of these labeling techniques, an alternative approach is for domain experts to craft rules or heuristics that can then be applied to unlabeled data [38]. These rules are called LFs and will be discussed in the next subsection.
2.3 Weak Supervision and LFs
@@
Weak supervision is a machine learning approach that addresses the challenge of obtaining large amounts of accurately labeled data [38, 39]. Instead of relying solely on expensive and time-consuming manual annotation, weak supervision leverages noisy, imprecise, or incomplete sources of information to generate training labels [39, 40]. Different forms of weak supervision exist [41], including:
@@
  • Heuristics: Rule-based labeling using domain knowledge [40] (e.g., labeling a query as ‘bug-related’ if it contains the word ‘error’).
@@
  • Distant Supervision: Using an external knowledge base or database to automatically generate labels [42] (e.g., labeling code comments based on the presence of specific API calls).
@@
  • Third-party Models: Using pre-trained models to generate labels for new data, even if the models were trained on different but related tasks [43].
@@
The key advantage of weak supervision is its ability to significantly reduce the time and cost associated with data labeling while enabling the use of larger datasets [38]. This can lead to the development of more robust and accurate models. Moreover, weak supervision, especially through the use of heuristics, allows for the direct integration of valuable domain expertise into the labeling process [26].
@@
LFs are a common and practical way to implement weak supervision. They are essentially a set of heuristic rules or predefined conditions to assign labels to data [38, 44, 27]. Instead of manually labeling each piece of data, a domain expert formulates a set of rules or conditions to create an LF. The primary goal of LFs is to facilitate the labeling process by determining the appropriate labels for each piece of data [38, 44, 27]. LFs offer several key advantages over manual labeling, including:
@@
  • Reusability: Once crafted, LFs can be applied multiple times across various batches of data [38, 44].
@@
  • Efficiency: LFs can significantly streamline the labeling process by automating label assignment based on well-defined criteria [38, 44, 27].
@@
Despite these advantages, developing effective LFs can be challenging, as it requires deep domain expertise and substantial time investment, especially as the complexity of the data and the rules increases [20]. To streamline the management and application of our LFs, we selected Snorkel as the underlying framework [27]. Snorkel is a framework specifically designed to enable and manage weak supervision through the use of LFs [38]. It is developed for programmatically building and managing training datasets without manual labeling [38]. Snorkel provides a structured approach for writing LFs in Python and uses a generative model to combine their outputs to create probabilistic labels for unlabeled data. Each LF in Snorkel is a standalone function that labels a subset of the data based on a certain rule or heuristic [38].
@@
Figure 3: Implementation of a Snorkel LF for identifying SE-related intents.
2.4 NLU-based Chatbots in the Era of LLMs
Large Language Models (LLMs) have transformed many fields, including SE [45, 46]. However, this raises the question of the relevance of NLU-based chatbots in the era of LLMs. LLMs and deep learning models require a substantial amount of labeled data for effective fine-tuning to be employed in intent classification tasks. Nevertheless, prior work has shown a scarcity of data in the SE domain [12], which hinders the utilization of deep learning models in labeling SE datasets. Even implementing the retrieval augmented generation using LLMs [47, 48, 49], may struggle to capture the nuances and specific terminology of SE tasks [50, 48]. While they can generate human-like responses, the accuracy and specificity of their outputs may vary, potentially lacking the precision required for SE context [50].
In contrast, NLU-based chatbots, which require training on smaller SE datasets compared to LLMs, excel in comprehending and addressing common development tasks, such as retrieving information about commits, branches, and issues [51, 4, 52]. By leveraging predefined intents and entities, these chatbots provide precise responses tailored to the SE context. Their deterministic nature enables developers to interpret and trace the reasoning behind each response, ensuring transparency and trust in the chatbot’s answers [47, 53]. Since the NLU-based chatbot depends on predefined intents and entities to answer questions, it lacks generalizability. In other words, the NLU-based chatbot does not answer questions that it is not trained on.
Still, we argue that both NLU-based and LLM-based chatbots have their own use cases. For instance, for a chatbot designed to answer specific questions about a software repository (e.g., identifying the fixing commit for a certain bug), an NLU-based chatbot would be a more suitable solution as it can be tailored to the specific project. Conversely, for a chatbot intended to address a broad spectrum of software development questions (e.g., learning best practices, fixing exceptions), an LLM-based chatbot would be a better choice since LLMs can be fine-tuned using general software development Q&A platforms (e.g., Stack Overflow).
3 Approach

### v2

@@
Abstract.
Software engineering (SE) chatbots are increasingly gaining attention for their role in enhancing development processes. At the core of chatbots are Natural Language Understanding platforms (NLUs), which enable them to comprehend user queries but require labeled data for training. However, acquiring such labeled data for SE chatbots is challenging due to the scarcity of high-quality datasets, as training requires specialized vocabulary and phrases not found in typical language datasets. Consequently, developers often resort to manually annotating user queries—a time-consuming and resource-intensive process. Previous approaches require human intervention to generate rules, called labeling functions (LFs), that categorize queries based on specific patterns. To address this issue, we propose an approach to automatically generate LFs by extracting patterns from labeled user queries. We evaluate our approach on four SE datasets and measure performance improvement from training NLUs on queries labeled by the generated LFs. The generated LFs effectively label data with AUC scores up to 85.3% and NLU performance improvements up to 27.2%. Furthermore, our results show that the number of LFs affects labeling performance. We believe that our approach can save time and resources in labeling users’ queries, allowing practitioners to focus on core chatbot functionalities rather than manually labeling queries.
@@ Abstract.
ccs: Computing methodologies Natural language processing
1. Introduction
@@
In the field of software engineering (SE), chatbots are deployed as conversational tools to automate a wide range of tasks, from assisting in problem-solving to answering questions about repositories (Grudin and Jacques, 2019; et al., 2023b, 2022). At the core of these chatbots are Natural Language Understanding platforms (NLUs). These NLUs serve as the backbone of the chatbot, responsible for interpreting human language into structured data that chatbots can act on (Abdellatif et al., 2021). They accomplish this by using machine learning and natural language processing techniques to dissect user queries, identifying the user’s intent and extracting key entities like bug IDs or version numbers (Prajwal et al., 2019). By doing so, NLUs make it possible for the chatbot to generate responses that are contextually relevant and specific to the user’s queries.
@@
The effectiveness of NLUs heavily relies on the availability of a large volume of high-quality training data (Zarcone et al., 2021; Lin et al., 2023; et al., 2021a). However, prior work shows that obtaining such data is expensive, especially in specialized domains like SE (Lin et al., 2020; Dominic et al., 2020; Abdellatif et al., 2020), where the domain-specific language and terminology pose unique challenges for data collection and labeling (Abdellatif et al., 2021). For example, common words like ‘push’, ‘fork’, and ‘commit’ have distinct meanings in the context of software development, which differ from their everyday usage (Abdellatif et al., 2021). This makes it difficult to rely on general-purpose datasets for training SE chatbots. Furthermore, there is a notable scarcity of publicly available, high-quality datasets for training chatbots in the SE domain (Motger et al., 2021). As a result, gathering sufficient and relevant training data for SE chatbots often requires significant effort and resources.
@@
To obtain the necessary training data for SE chatbots, practitioners often resort to mining chatbot user queries as a data source (Farhour et al., 2024). However, chatbot developers need to label these queries before they can be used for training the NLU, introducing significant challenges (et al., 2017b). First, manual labeling is labor-intensive and can incur substantial costs (Tu et al., 2020), particularly when domain expertise is essential for ensuring the accuracy of labels (Pinhanez, 2019; Zhang et al., 2016). This increases the financial burden and extends the time required to prepare data for training purposes (Ghahreman and Dastjerdi, 2011; Tu et al., 2020; Pujara et al., 2011). Second, although semi-automated labeling methods exist, they still require human intervention, such as domain experts developing heuristics for labeling user queries (Galhotra et al., 2021; et al., 2021b, 2018b). These heuristics are also called Labeling Functions (referred hereafter simply as LFs). Specifically, LFs are programmatic rules or functions that assign labels to data points based on certain conditions or patterns. For example, an LF for an SE chatbot might label a query as ‘bug-related’ if it contains keywords like ‘error’, ‘issue’, or ‘fix’.
@@
In response to the challenges outlined above regarding data labeling, we introduce our approach that automates the process of generating LFs specifically for SE datasets. To accomplish this, our approach takes a small set of labeled data as input and automatically analyzes and extracts patterns from it, and uses those patterns to generate the LFs capable of auto-labeling data. After the generation of LFs, we use them to label data and train the NLU of the chatbot on the trained data (Bocklisch et al., 2017). Our approach is organized into three main components. First, the Grouper is responsible for expanding the initial labeled data by identifying similar queries (Schopf et al., 2023). Second, the LF Generator takes on the role of extracting patterns from this expanded data to create LFs (Bringer et al., 2019). Finally, the Pruner filters out low-quality LFs from the pool of generated LFs (Varma and Ré, 2018; et al., 2020a). We evaluate our approach based on four datasets (namely, AskGit, MSA, Ask Ubuntu, and Stack Overflow) used to develop chatbots for performing various SE tasks. Specifically, we aim to answer the following research questions in this paper:
@@
    How well do the generated LFs label data? We evaluate the quality of the generated LFs in terms of their effectiveness in labeling data, and subsequently, in training SE chatbots (Yang et al., 2022; Ferri et al., 2009). Our analysis shows that the generated LFs effectively label data with AUC scores of above 75.5% for three out of the four studied datasets (except Stack Overflow). Additionally, we observe that for these three datasets, using the auto-labeled data for training can enhance the NLU’s performance, with AUC score improvements of up to 27.2%.
@@
  • We make our dataset and code publicly available to facilitate future research in this area (E. Alor and Shihab, 2024).
Paper Organization. The remainder of the paper is organized as follows. Section 2 provides the background that forms the basis for our study. Section 3 presents the details of our approach and its key components. Section 4 describes the setup for our empirical study followed by Section 5 which presents the results of our RQs. Section 6 discusses the relevance of NLU-based chatbots in the era of LLMs and analyzes the impact of the number of LFs on labeling performance. Section 7 discusses threats to the validity of our study. Finally, Section 8 reviews the related work and Section 9 concludes the paper.
2. Background
Before proceeding to our approach, we explain in this section the chatbot-related terminologies, data labeling process, and LFs. Also, we discuss the role of NLU-based chatbots in the era of LLMs.
2.1. Chatbot Training and Data Challenges
Software chatbots serve as the conduit between users and services (Lebeuf et al., 2017). Users input their queries to the chatbot in natural language, which then performs the requested action (e.g., querying the database) and responds to the user’s question. NLU platforms are the backbone for chatbots to understand the user’s question. Specifically, NLU extracts two key aspects from the input query: the topic of the user’s query or what is being asked about (known as the ‘Entity’) and what the user is asking it to do (known as the ‘Intent’).
Similar to any machine learning model, NLUs need to be trained to extract intents and entities. In particular, for each intent, the NLU needs to be trained on a set of queries that represent different ways a user could express that intent. Therefore, the training data should include a wide range of examples for each intent to capture the variety of ways in which people express their needs. Figure 1 presents a snapshot of training data for the FileCreator and IssueClosingDate intents with their training examples. Chatbot practitioners need to brainstorm the different ways that the user could ask the question for a specific intent (Abdellatif et al., 2020) to train the NLU. Nevertheless, this is a resource-intensive and time-consuming task (Farhour et al., 2024).
@@ 2.1. Chatbot Training and Data Challenges
Figure 1. Example dataset for training an SE chatbot.
Alternatively, chatbot practitioners leverage user-chatbot conversations to augment their dataset (Farhour et al., 2024). To demonstrate this, we present the user-chatbot interaction example shown in Figure 2. In the example, the user (highlighted in green) asks the chatbot software repository-related questions. For the first query, the user asks, ‘I want to know who created map.json’. Here, the intent is to find out the developer who created the file. The chatbot’s NLU analyzes this query, extracting the intent (FileCreator with a confidence score of 0.87) and the relevant entity (file_name is map.json). Based on this understanding, the chatbot then provides a response identifying the file creator.
Figure 2. Example of a user’s interaction with an SE chatbot.
@@ 2.2. Data Labeling
The process of data labeling involves assigning appropriate labels to a dataset (et al., 2020a; Galhotra et al., 2021). There are several data labeling techniques that are traditionally used (Prenner and Robbes, 2021; Robbes and Janes, 2019). These include manual labeling, crowd-sourcing, and hiring domain experts (Prenner and Robbes, 2021; Tu et al., 2020; Pinhanez, 2019; Rodríguez and Müller, 2012).
@@
Manual labeling is the annotation of unlabeled data by humans (et al., 2017b). The process typically begins with gathering the unlabeled data (Hancock et al., 2019) and reviewing them. Next, appropriate labels for intents and entities are assigned to the unlabeled data, categorizing them into classes (Hancock et al., 2019). Once the data is suitably annotated, it is incorporated into the chatbot’s training dataset to further enhance its accuracy (Hancock et al., 2019). Because of all these steps, the manual labeling process requires significant effort and is time-consuming (Tu et al., 2020). It can also lead to inconsistencies due to human error, making it less practical for large-scale projects (et al., 2021a).
@@
Crowdsourcing involves distributing the task of labeling data to a large group of people, often through an online platform. This method can significantly accelerate the data labeling process by leveraging the crowd’s collective effort. Although crowdsourcing can help label a large volume of data quickly, it may not always guarantee the quality necessary for specialized domains like SE. The lack of domain-specific knowledge among the crowd can lead to inaccuracies in labeling, affecting the overall quality of the training data (Zhang et al., 2016; Sheng and Zhang, 2019; Stolee and Elbaum, 2010).
@@
Recruiting domain experts refers to engaging individuals with specialized knowledge in a particular area, such as SE, to label the data. These experts bring a high level of accuracy and insight to the labeling process, ensuring the data is correctly annotated with the appropriate intents and entities. However, this method can be costly and may not scale well for large datasets, making it a challenging option for projects with limited budgets (Pinhanez, 2019; Zhang et al., 2016). Due to the limitations of these labeling techniques, an alternative approach is for domain experts to craft rules or heuristics that can then be applied to unlabeled data (et al., 2017a). These rules are called Labeling Functions (LFs) and will be discussed in the next subsection.
2.3. Weak Supervision and LFs
@@
Weak supervision is a machine learning approach that addresses the challenge of obtaining large amounts of accurately labeled data (et al., 2017a; Karamanolakis et al., 2021). Instead of relying solely on expensive and time-consuming manual annotation, weak supervision leverages noisy, imprecise, or incomplete sources of information to generate training labels (Karamanolakis et al., 2021; Boecking et al., 2020). Different forms of weak supervision exist (Hernández-González et al., 2016), including:
@@
  • Heuristics: Rule-based labeling using domain knowledge (Boecking et al., 2020) (e.g., labeling a query as ‘bug-related’ if it contains the word ‘error’).
@@
  • Distant Supervision: Using an external knowledge base or database to automatically generate labels (Mintz et al., 2009) (e.g., labeling code comments based on the presence of specific API calls).
@@
  • Third-party Models: Using pre-trained models to generate labels for new data, even if the models were trained on different but related tasks (Mann and McCallum, 2010).
@@
The key advantage of weak supervision is its ability to significantly reduce the time and cost associated with data labeling while enabling the use of larger datasets (et al., 2017a). This can lead to the development of more robust and accurate models. Moreover, weak supervision, especially through the use of heuristics, allows for the direct integration of valuable domain expertise into the labeling process (Varma and Ré, 2018).
@@
LFs are a common and practical way to implement weak supervision. They are essentially a set of heuristic rules or predefined conditions to assign labels to data (et al., 2017a; Lison et al., 2021; et al., 2020a). Instead of manually labeling each piece of data, a domain expert formulates a set of rules or conditions to create an LF. The primary goal of LFs is to facilitate the labeling process by determining the appropriate labels for each piece of data (et al., 2017a; Lison et al., 2021; et al., 2020a). LFs offer several key advantages over manual labeling, including:
@@
  • Reusability: Once crafted, LFs can be applied multiple times across various batches of data (et al., 2017a; Lison et al., 2021).
@@
  • Efficiency: LFs can significantly streamline the labeling process by automating label assignment based on well-defined criteria (et al., 2017a; Lison et al., 2021; et al., 2020a).
@@
Despite these advantages, developing effective LFs can be challenging, as it requires deep domain expertise and substantial time investment, especially as the complexity of the data and the rules increases (Galhotra et al., 2021). To streamline the management and application of our LFs, we selected Snorkel as the underlying framework (et al., 2020a). Snorkel is a framework specifically designed to enable and manage weak supervision through the use of LFs (et al., 2017a). It is developed for programmatically building and managing training datasets without manual labeling (et al., 2017a). Snorkel provides a structured approach for writing LFs in Python and uses a generative model to combine their outputs to create probabilistic labels for unlabeled data. Each LF in Snorkel is a standalone function that labels a subset of the data based on a certain rule or heuristic (et al., 2017a).
@@
Figure 3. Implementation of a Snorkel LF for identifying SE-related intents.
2.3.1. Benefits of an LF Generation Approach
Given the capabilities of LLMs, one might ask why our approach for generating the LFs themselves does not primarily rely on LLMs (e.g., by prompting an LLM to write labeling rules). Our methodology, using semantic similarity (Grouper), pattern/statistic extraction and ML classifiers (Generator), and quality filtering (Pruner), was a deliberate choice based on several factors prioritizing transparency, control, efficiency, and alignment with established weak supervision principles:
  • SE Data Scarcity for LLM Fine-Tuning: LLMs typically require substantial amounts of domain-specific labeled data for effective fine-tuning (Shin et al., 2023), yet such annotated datasets are particularly scarce in specialized SE contexts. Our approach addresses this fundamental challenge by automatically generating LFs that can create labeled data from readily available unlabeled queries, effectively bootstrapping the data annotation process without relying on pre-existing large-scale SE datasets.
  • Reusability and Computational Cost: A key output of our approach is a set of explicit LFs. These LFs are reusable artifacts; once generated, they can be applied consistently to label future batches of unlabeled data as they arrive (et al., 2017a; Lison et al., 2021). This contrasts with using an LLM for direct, one-time classification of a dataset, where the underlying labeling logic might not be captured as an easily reusable function for ongoing labeling needs. Reusability helps to save cost and energy resources. Also our approach utilizes relatively lightweight tools like Sentence Transformers, CountVectorizer, and standard ML libraries (e.g., scikit-learn), which are computationally less demanding and more accessible than requiring large LLMs for the core pattern-finding and rule-generation steps themselves (Patterson et al., 2021).
  • Explainability and Trustworthiness: Our generated LFs are based on identifiable features (unique words, entity types, co-occurrences) or explainable ML models. This allows practitioners to inspect why a query receives a certain label, fostering trust and enabling easier debugging. In contrast, LLM-generated rules often function as “black boxes” that make verification difficult and may produce inconsistent labels for similar queries.
In essence, we traded the potential generative power of LLMs in the LF creation step for greater transparency, control, and efficiency in the labeling process itself, aligning with the practical needs of building reliable data labeling pipelines for specialized domains like SE.
3. Approach